If there are two whole numbers that when added the sum is 118, and when the smaller is subtracted from the larger the difference is 36; What are the numbers?
Answer by Earlsdon(6294)

Let a = the larger number and b = the smaller number.
a+b = 118
a-b = 36 Rewrite this as a = b+36 and substitute into the first equation.
(b+36)+b = 118 Simplify and solve for b.
2b+36 = 118 Subtract 36 from both sides.
2b = 82 Divide both sides by 2.
b = 41 and...
a = b+36
a = 41+36
a = 77